OpenStudy (anonymous):

determine whether each of the ff is a linear equation. 1. y=-5 2. x-2y=4

OpenStudy (anonymous):

is not x is there in eq 1

OpenStudy (anonymous):

x-2y=4 -2y=-x+4 2y=x+4 y=x/2+2 so it is a linear eq

OpenStudy (anonymous):

for first It's the equation of a horizontal line going through the y axis at -5; therefore it's a linear equation.